OpenStudy (anonymous):

Evaluate the following expression. (a + b)^0

OpenStudy (shamil98):

(a+b)^ 0 = 1

OpenStudy (anonymous):

anything to power zero ( except zero) is 1

OpenStudy (shamil98):

anything to the power of 0 is 1, with the exception of 0.

OpenStudy (anonymous):

(a+b)^ 0 = 1

OpenStudy (anonymous):

anything to the power of 0 is 1, with the exception of 0.

OpenStudy (anonymous):

what about this 2 ^-2

OpenStudy (shamil98):

2^-2 = 1/2^2 1/4 = 0.25

OpenStudy (anonymous):

Rewrite using a negative exponent. 1/3^2

OpenStudy (shamil98):

3^ -2

OpenStudy (anonymous):

ok last one 1/(xy)^+1

OpenStudy (anonymous):

@magan , did it in the other window :D

OpenStudy (anonymous):

yes thank you